# Stochastic Re-weighted Gradient Descent via Distributionally Robust Optimization

### v4

@@ 4.3 Supervised Learning
ERM (Gulrajani & Lopez-Paz, 2020) 85.5 $\pm$ 0.1 77.5 $\pm$ 0.4 66.5 $\pm$ 0.2 40.9 $\pm$ 0.1 67.6
MIRO (Cha et al., 2022) 85.4 $\pm$ 0.4 79.0 $\pm$ 0.0 70.5 $\pm$ 0.4 44.3 $\pm$ 0.2 69.8
@@
4.4 Tabular Classification
Learning with tabular data is a task where traditional machine learning methods, like random forest (Breiman, 2001; Friedman, 2001) are incredibly competitive against deep learning-based methods (Yoon et al., 2020). Recently, Majmundar et al. (2022) obtained SOTA results for tabular classification using self-supervised representation learning and relying on the learned representations in the downstream classification tasks (see Section 1). Their work developed two algorithms namely, MET (representation learning with adversarial training) and MET-S (representation learning without adversarial training). The adversarial training adds robustness to the learned representations, thus improving performance. In this experiment, we integrate RGD with MET-S instead of doing adversarial training. This allows us to test the robustness properties of the models trained with RGD. Table 6 and Table 5 shows gains on multiple tabular datasets for the multi-class classification and binary classification tasks. Notably, our approach outperforms previous SOTA in this problem by +1.27%, and +1.5% on the multi-class and binary classification tasks respectively. We refer to Appendix F.6 for a comprehensive comparison with baselines such as Gradient Boosting Decision Trees (Friedman, 2001), VIME (Yoon et al., 2020), SubTab (Ucar et al., 2021), TabNet (Arik & Pfister, 2019), DACL+ (Verma et al., 2021) and many more. Our motivation to experiment on these “permuted” MNIST, “permuted” CIFAR, and “permuted” FMNIST can be traced back to the introduction of these datasets in the works of (Yoon et al., 2020; Ucar et al., 2021). Subsequently, other recent works such as (Majmundar et al., 2022) also experimented on these datasets and have become a standard benchmark for tabular classification.
4.5 Out Of Domain Generalization
In this section, we show that our technique can be used to boost the performance of OOD generalization techniques. We experiment on DomainBed, a standard benchmark used to study the out-of-domain performance of models. More information about the benchmark, the task to solve, and the metric is discussed in Appendix C. The benchmark is notorious since the most basic approach, such as straightforward Empirical Risk Minimization (ERM) as evaluated by (Gulrajani & Lopez-Paz, 2020), was the SOTA method for a long time. Most new approaches either performed worse than ERM or marginally better. In recent years, breakthroughs such as MIRO (Cha et al., 2022) and FRR (Addepalli et al., 2022) have pushed the problem further by significantly improving the benchmarks. We show that integrating our proposed approach RGD with these approaches (specifically FRR) significantly improves performance (an average of +0.7%). Table 7 illustrates the accuracy performance numbers of a few baseline methods and our proposed approach. A more comprehensive comparison with additional baselines such as IRM (Arjovsky et al., 2019), CORAL (Sun & Saenko, 2016), MTL (Blanchard et al., 2021), SagNet (Nam et al., 2021), and many more in the Table 22. We further depict the environment-wise breakdown of the accuracy of each of the baseline algorithms in Appendix F.8.
4.6 Meta-Learning
In meta-learning, where the objective is to efficiently learn new tasks with limited data, the re-weighing of tasks becomes crucial. While Empirical Risk Minimization (ERM), may perform well on common tasks, it can fail catastrophically on rare and challenging ones. Recently, (Li et al., 2023) tried KL-DRO in this domain. However, the authors applied their algorithm for solving meta-regression on a toy-dataset which has no outliers. Whereas, we apply our algorithm on practically relevant and real-world datasets. Building upon the experimental results of Kumar et al. (2022), we make comparisons with our MAML + RGD approach as the proposed variant. We evaluate RGD not only based on the average performance across tasks but also on the Worst-Kof tasks in a fixed task pool. Our experiments on various benchmarks, including Omniglot 5-way 1-shot, Omniglot 20-way 1-shot, and miniImageNet 5-way 1-shot, demonstrate significant improvements in the Worst-Kmetric (Table 8). For example, on Omniglot 20-way 1-shot, our proposed reweighting scheme improves overall performance by 1.83% and the Worst-10% performance by 2.28%. Similarly, on the challenging miniImageNet 5-way 1-shot benchmark, we achieve a substantial improvement of approximately 3% across the board. Further results in this domain are discussed in Appendix F.7.
5 Conclusion and Future Work
@@
One potential avenue for future research is to theoretically understand the weight clipping in our algorithm. While weight clipping can help us withstand benign outliers, we believe it is not the best way to handle a large number of systematic or adversarial corruptions. In the future, we plan to develop variants of RGD that can tolerate adversarial corruptions in the training data, while simultaneously improving the model generalization. Additionally, we plan to evaluate our technique on large-scale tasks, such as fine-tuning Large Language Models (LLMs) and other foundation models. This will help us better understand the usefulness and limitations of our approach.
@@
Acknowledgements
We would like to thank Ahmad Beirami, Virginia Smith, Manzil Zaheer, Tian Li, and Maziar Sanjabi for providing detailed feedback on our paper. We would like to thank Elan Rosenfeld, and Tianbao Yang for pointing us to important prior works. We extend our sincere gratitude to Prateek Jain, Pradeep Shenoy, Anshul Nasery, Lovish Madaan, and the numerous dedicated members of the machine learning and optimization team at Google Research India for their invaluable feedback and contributions to this work.
@@ A.1 Per-Sample Reweighting
In this section, we review data reweighting techniques developed outside of the DRO community. The idea of re-weighting samples can be dated back to the works of (Chawla et al., 2002; Zadrozny, 2004), which pre-computed per-sample weights using certain prior knowledge. Recent approaches alleviate the need for human supervision by dynamically computing the per-sample weights. One of the early works in this category is AdaBoost, which is a popular boosting algorithm (Freund & Schapire, 1997). Similar to RGD, AdaBoost uses exponential weighting mechanism to reweight data points. However, AdaBoost is used for learning an ensemble of weak learners. Whereas, in this work, we are interested in learning a single model that can achieve better generalization guarantees. Furthermore, AdaBoost is only studied for supervised learning (in particular, classification and regression). In contrast, RGD can be applied on any learning task. Recent works of (Leng et al., 2022; Lin et al., 2017) showed that certain modifications to standard cross entropy loss - that involve truncating its Taylor-series expansion - can improve the performance of DNNs. These techniques can be viewed as performing sample re-weighting. However, these techniques only apply to cross-entropy loss and are not easily extendable to general learning tasks.
@@
Other approaches based on meta-learning have been proposed for class imbalance and label noise (Shu et al., 2019; Ren et al., 2018; Gonzalez & Miikkulainen, 2021). Many popular approaches in this line of work require training a separate neural network for re-weighting the data points (Ren et al., 2018; Shu et al., 2019). However, these approaches are seldom used in practice as the underlying optimization bi-level optimization problem is hard to implement (Ivanova & Ablin, 2023). Unlike these approaches, our RGD algorithm does not require a separate neural network for re-weighting and thus doesn’t add any computational overhead over vanilla training. Moreover, compared to existing sample re-weighting techniques, our approach applies to various learning tasks (see Section 4). Another line of work uses a history buffer which stores a snapshot of the trajectory of each point and facilitates giving more importance to points which leads to more learning in the model (Zhang & Pfister, 2021). Other approaches, such as (Zhu et al., ), use reinforcement learning to learn the per-sample weights using a “pretraining-boosting” two-stage MDP curriculum where the agent network is firstly pre-trained and optimized for deployment in the classification problem. Another line of work has considered sample re-weighting in the presence of outliers (Kumar et al., 2010; De La Torre & Black, 2003; Jiang et al., 2014a, b; Wang et al., 2017; Li et al., 2020, 2023). These works down-weight points with high-loss value. The rationality behind this lies in the idea that these high-loss samples are more likely to be outliers and, thus, should be ignored during the training process. Finally, works such (Castells et al., 2020) propose a confidence-aware loss proportional to the lowest loss of that sample. They use a threshold ($\gamma$) to decide how practical or important each point is.
@@ A.2 Pre-conditioning
Pre-conditioning can usually mean normalization of inputs, batch normalization, or scaling gradients in a few directions. This section predominantly discusses techniques that focus on scaling gradients in a few directions. A common technique to improve the training speed in deep learning is using adaptive step-size optimization methods, such as the Newton method, which takes advantage of the second-order gradients. However, computing the Hessian matrix is computationally intensive, leading to Quasi-Newton methods: methods that approximate the value of the Hessian instead of computing them every time (Le et al., 2011). Another popular alternative is to use an element-wise adaptive learning rate, which has shown great promise in deep learning. Some of the popular techniques here include ADAgrad (Duchi et al., 2011), RMSProp (Ruder, 2016), ADAdelta (Zeiler, 2012). For instance, ADAgrad is a diagonal pre-conditioning technique where the pre-conditioning across each dimension is computed as the inverse square root of the norms of gradients along that dimension accumulated over training. Unfortunately, this accumulation of gradients makes it susceptible to falling in a saddle point as the scaling factor decreases monotonically.
@@ A.3 Curriculum Learning
Another important research area that has explored data reweighting is Curriculum Learning (CL). CL, originally introduced by (Bengio et al., 2009), is a vast domain focussing on how the model should be taught, and draws inspiration from how humans learn concepts. For instance, humans generally grasp on to easier concepts such as basic shapes (triangle, rectangle, etc.) before moving on to learning significantly more complex structures (heptagram, triquetra, etc.). Curriculum learning strategies have been widely used in various areas of machine learning and involves finding a way to rank samples, as well as the right pacing functions for introducing more difficult data in our training. The techniques developed for CL have typically focused on giving importance to easier samples at the beginning of training, and slowly progressing towards harder samples (Bengio et al., 2009; Chen & Gupta, 2015; Tudor Ionescu et al., 2016; Pentina et al., 2015; Shi et al., 2015; Spitkovsky et al., 2009; Zaremba & Sutskever, 2014). In contrast, DRO focuses the learning on harder samples throughout the training process. That being said, there have also been a class of works in CL which showed the learning harder examples first, and then moving to easier ones could lead to improved performance in certain conditions, through Hard Example mining (HEM) or anti-curriculum (Jesson et al., 2017; Shrivastava et al., 2016; Wang & Vasconcelos, 2018; Zhou et al., 2020; Braun et al., 2017; Pi et al., 2016). There have been predominantly three classes of CL, and various amalgamations of these in literature (Soviany et al., 2022).
Vanilla CL: The vanilla CL usually involves a pre-defined notion of hardness. For example, (Bengio et al., 2009) used geometric shapes to clearly differentiate easy and hard samples. Others such as (Spitkovsky et al., 2009) exploited the length of sequences as a signal for difficulty.
Self-Paced Learning (SPCL): This differs from the vanilla CL with respect to the evaluation of difficulty. This concept of “difficulty” is not known beforehand and is measured repeatedly during training. Works such as (Kumar et al., 2010) used the likelihood of the prediction to rank the samples. Other works such as (Lee & Grauman, 2011) used objectness as a measure to define the training schedule.
@@
In this work, we will resort to only describing few works which focus on instance level reweighting of data points (Kumar et al., 2010; Li et al., 2017; Kumar et al., 2011; Pi et al., 2016; Liang et al., 2016; Fan et al., 2017; Li et al., 2017). Most of these works (Kumar et al., 2011; Fan et al., 2017) follow a binary weighting mechanism of $\{0,1\}$ to decide whether the model should learn using the current sample or not. Others such as (Li et al., 2017; Pi et al., 2016; Liang et al., 2016) are more continuous in the weighing mechanism and generally give higher weights to samples with lower losses. This makes them fundamentally different from RGD attempts to achieve in this work. RGD attempts to focus more heavily on the harder samples throughout training and does so using a simple closed form expression of the loss.
@@ Appendix B Proofs of Section 3
[…]
@@
For any fixed $\beta,\theta$, the inner supremum is attained at a $P^{\prime}$ that satisfies (see Theorem 1 of (Hsieh et al., 2019))

### v5

@@ 4.4 Tabular Classification
Learning with tabular data is a task where traditional machine learning methods, like random forest Breiman (2001); Friedman (2001) are incredibly competitive against deep learning-based methods (Yoon et al., 2020). Recently, Majmundar et al. (2022) obtained SOTA results for tabular classification using self-supervised representation learning and relying on the learned representations in the downstream classification tasks (see Section 1). Their work developed two algorithms namely, MET (representation learning with adversarial training) and MET-S (representation learning without adversarial training). The adversarial training adds robustness to the learned representations, thus improving performance. In this experiment, we integrate RGD with MET-S instead of doing adversarial training. This allows us to test the robustness properties of the models trained with RGD. Table 7 and Table 6 shows gains on multiple tabular datasets for the multi-class classification and binary classification tasks. Notably, our approach outperforms previous SOTA in this problem by +1.27%, and +1.5% on the multi-class and binary classification tasks respectively. We refer to Appendix D.6 for a comprehensive comparison with baselines such as Gradient Boosting Decision Trees (Friedman, 2001), VIME (Yoon et al., 2020), SubTab (Ucar et al., 2021), TabNet (Arik & Pfister, 2021), DACL+ (Verma et al., 2021) and many more. Our motivation to experiment on these “permuted” MNIST, “permuted” CIFAR, and “permuted” FMNIST can be traced back to the introduction of these datasets in the works of Yoon et al. (2020); Ucar et al. (2021). Subsequently, other recent works such as Majmundar et al. (2022) also experimented on these datasets and have become a standard benchmark for tabular classification.
4.5 Out Of Domain Generalization
@@
ERM Gulrajani & Lopez-Paz (2021) 85.5 $\pm$ 0.1 77.5 $\pm$ 0.4 66.5 $\pm$ 0.2 40.9 $\pm$ 0.1 67.6
MIRO Cha et al. (2022) 85.4 $\pm$ 0.4 79.0 $\pm$ 0.0 70.5 $\pm$ 0.4 44.3 $\pm$ 0.2 69.8
@@
In this section, we show that our technique can be used to boost the performance of OOD generalization techniques. We experiment on DomainBed, a standard benchmark used to study the out-of-domain performance of models. More information about the benchmark, the task to solve, and the metric is discussed in Appendix D.8.1. The benchmark is notorious since the most basic approach, such as straightforward Empirical Risk Minimization (ERM) as evaluated by Gulrajani & Lopez-Paz (2021), was the SOTA method for a long time. Most new approaches either performed worse than ERM or marginally better. In recent years, breakthroughs such as MIRO (Cha et al., 2022) and FRR (Addepalli et al., 2023) have pushed the problem further by significantly improving the benchmarks. We show that integrating our proposed approach RGD with these approaches (specifically FRR) significantly improves performance (an average of +0.7%). Table 8 illustrates the accuracy performance numbers of a few baseline methods and our proposed approach. A more comprehensive comparison with additional baselines such as IRM (Arjovsky et al., 2019), CORAL (Sun & Saenko, 2016), MTL (Blanchard et al., 2021), SagNet (Nam et al., 2021), and many more in the Table 24. We further depict the environment-wise breakdown of the accuracy of each of the baseline algorithms in Appendix D.8.
4.6 Meta-Learning
@@
In meta-learning, the goal is to learn representations that generalize effectively to new tasks, even when provided with limited examples. However, task heterogeneity poses a significant challenge. Some tasks may be inherently simpler to learn, leading models to prioritize these and neglect the more difficult, less frequent tasks. While Empirical Risk Minimization (ERM) may perform well on common tasks, its performance can deteriorate drastically on rare and challenging ones. This necessitates a mechanism for re-weighting tasks to ensure balanced learning. Building upon the experimental results of Kumar et al. (2023), we make comparisons with our MAML + RGD approach as the proposed variant. We evaluate RGD not only based on the average performance across tasks but also on the Worst-Kof tasks in a fixed task pool. Our experiments on various benchmarks, including Omniglot 5-way 1-shot, Omniglot 20-way 1-shot, and miniImageNet 5-way 1-shot, demonstrate significant improvements in the Worst-Kmetric (Table 9). For example, on Omniglot 20-way 1-shot, our proposed reweighting scheme improves overall performance by 1.83% and the Worst-10% performance by 2.28%. Similarly, on the challenging miniImageNet 5-way 1-shot benchmark, we achieve a substantial improvement of approximately 3% across the board. Further results in this domain are discussed in Appendix D.7.
5 Conclusion, Limitations and Future Work
@@
While RGD shows promising results, it has the following limitations that warrant further investigation: (a) outlier handling: our current approach uses weight clipping to mitigate the impact of outliers. While empirically effective, a more principled approach to outlier robustness in DRO is worth exploring, and (b) robustness to noise: while RGD can handle benign noise, it can fail in the presence of adversarial/systematic noise. In the future, we plan to develop variants of RGD that can tolerate adversarial corruptions in the training data, while simultaneously improving the model generalization. Additionally, we plan to evaluate our technique on large-scale tasks, such as fine-tuning Large Language Models (LLMs) and other foundation models. This will help us better understand the usefulness and limitations of our approach.
@@
Acknowledgments
We would like to thank Ahmad Beirami, Virginia Smith, Manzil Zaheer, Tian Li, and Maziar Sanjabi for providing detailed feedback on our paper. We would like to thank Elan Rosenfeld, and Tianbao Yang for pointing us to important prior works. We extend our sincere gratitude to Prateek Jain, Pradeep Shenoy, Anshul Nasery, Lovish Madaan, and the numerous dedicated members of the machine learning and optimization team at Google DeepMind India for their invaluable feedback and contributions to this work.
@@ A.1 Per-Sample Reweighting
In this section, we review data reweighting techniques developed outside of the DRO community. The idea of re-weighting samples can be dated back to the works of Chawla et al. (2002); Zadrozny (2004), which pre-computed per-sample weights using certain prior knowledge. Recent approaches alleviate the need for human supervision by dynamically computing the per-sample weights. One of the early works in this category is AdaBoost, which is a popular boosting algorithm (Freund & Schapire, 1997). Similar to RGD, AdaBoost uses exponential weighting mechanism to reweight data points. However, AdaBoost is used for learning an ensemble of weak learners. Whereas, in this work, we are interested in learning a single model that can achieve better generalization guarantees. Furthermore, AdaBoost is only studied for supervised learning (in particular, classification and regression). In contrast, RGD can be applied on any learning task. Recent works of Leng et al. (2022); Lin et al. (2017) showed that certain modifications to standard cross entropy loss - that involve truncating its Taylor-series expansion - can improve the performance of DNNs. These techniques can be viewed as performing sample re-weighting. However, these techniques only apply to cross-entropy loss and are not easily extendable to general learning tasks.
@@
Other approaches based on meta-learning have been proposed for class imbalance and label noise (Shu et al., 2019; Ren et al., 2018; Gonzalez & Miikkulainen, 2021). Many popular approaches in this line of work require training a separate neural network for re-weighting the data points (Ren et al., 2018; Shu et al., 2019). However, these approaches are seldom used in practice as the underlying bi-level optimization problem is hard to implement (Ivanova & Ablin, 2023). Unlike these approaches, our RGD algorithm does not require a separate neural network for re-weighting and thus doesn’t add any computational overhead over vanilla training. Moreover, compared to existing sample re-weighting techniques, our approach applies to various learning tasks (see Section 4). Another line of work uses a history buffer which stores a snapshot of the trajectory of each point and facilitates giving more importance to points which leads to more learning in the model (Zhang & Pfister, 2021). Other approaches, such as Zhu et al. (2022b), use reinforcement learning to learn the per-sample weights using a “pretraining-boosting” two-stage MDP curriculum where the agent network is firstly pre-trained and optimized for deployment in the classification problem. Another line of work has considered sample re-weighting in the presence of outliers (Kumar et al., 2010; De La Torre & Black, 2003; Jiang et al., 2014a; b; Wang et al., 2017; Li et al., 2021; 2023). These works down-weight points with high-loss value. The rationality behind this lies in the idea that these high-loss samples are more likely to be outliers and, thus, should be ignored during the training process. Finally, works such Castells et al. (2020) propose a confidence-aware loss proportional to the lowest loss of that sample. They use a threshold ($\gamma$) to decide how practical or important each point is.
@@ A.2 Pre-conditioning
Pre-conditioning can usually mean normalization of inputs, batch normalization, or scaling gradients in a few directions. This section predominantly discusses techniques that focus on scaling gradients in a few directions. A common technique to improve the training speed in deep learning is using adaptive step-size optimization methods, such as the Newton method, which takes advantage of the second-order gradients. However, computing the Hessian matrix is computationally intensive, leading to Quasi-Newton methods: methods that approximate the value of the Hessian instead of computing them every time Le et al. (2011). Another popular alternative is to use an element-wise adaptive learning rate, which has shown great promise in deep learning. Some of the popular techniques here include ADAgrad (Duchi et al., 2011), RMSProp (Ruder, 2016), ADAdelta (Zeiler, 2012). For instance, ADAgrad is a diagonal pre-conditioning technique where the pre-conditioning across each dimension is computed as the inverse square root of the norms of gradients along that dimension accumulated over training. Unfortunately, this accumulation of gradients makes it susceptible to falling in a saddle point as the scaling factor decreases monotonically.
@@ A.3 Curriculum Learning
Another important research area that has explored data reweighting is Curriculum Learning (CL). CL, originally introduced by Bengio et al. (2009), is a vast domain focussing on how the model should be taught, and draws inspiration from how humans learn concepts. For instance, humans generally grasp on to easier concepts such as basic shapes (triangle, rectangle, etc.) before moving on to learning significantly more complex structures (heptagram, triquetra, etc.). Curriculum learning strategies have been widely used in various areas of machine learning and involves finding a way to rank samples, as well as the right pacing functions for introducing more difficult data in our training. The techniques developed for CL have typically focused on giving importance to easier samples at the beginning of training, and slowly progressing towards harder samples (Bengio et al., 2009; Chen & Gupta, 2015; Tudor Ionescu et al., 2016; Pentina et al., 2015; Shi et al., 2015; Spitkovsky et al., 2010; Zaremba & Sutskever, 2014). In contrast, DRO focuses the learning on harder samples throughout the training process. That being said, there have also been a class of works in CL which showed the learning harder examples first, and then moving to easier ones could lead to improved performance in certain conditions, through Hard Example mining (HEM) or anti-curriculum (Jesson et al., 2017; Shrivastava et al., 2016; Wang & Vasconcelos, 2018; Zhou et al., 2020; Braun et al., 2017; Pi et al., 2016). There have been predominantly three classes of CL, and various amalgamations of these in literature (Soviany et al., 2022).
Vanilla CL: The vanilla CL usually involves a pre-defined notion of hardness. For example, Bengio et al. (2009) used geometric shapes to clearly differentiate easy and hard samples. Others such as Spitkovsky et al. (2010) exploited the length of sequences as a signal for difficulty.
Self-Paced Learning (SPCL): This differs from the vanilla CL with respect to the evaluation of difficulty. This concept of “difficulty” is not known beforehand and is measured repeatedly during training. Works such as Kumar et al. (2010) used the likelihood of the prediction to rank the samples. Other works such as Lee & Grauman (2011) used objectness as a measure to define the training schedule.
@@
In this work, we will resort to only describing few works which focus on instance level reweighting of data points (Kumar et al., 2010; Li et al., 2017; Kumar et al., 2011; Pi et al., 2016; Liang et al., 2016; Fan et al., 2017; Li et al., 2017). Most of these works (Kumar et al., 2011; Fan et al., 2017) follow a binary weighting mechanism of $\{0,1\}$ to decide whether the model should learn using the current sample or not. Others such as Li et al. (2017); Pi et al. (2016); Liang et al. (2016) are more continuous in the weighing mechanism and generally give higher weights to samples with lower losses. This makes them fundamentally different from RGD attempts to achieve in this work. RGD attempts to focus more heavily on the harder samples throughout training and does so using a simple closed form expression of the loss.
A.4 Comparison with existing KL-DRO optimization approaches
In this section, we discuss a few more aspects of the related work that weren’t discussed in the main paper. We specifically focus on prior works that developed algorithms for KL divergence based DRO. The earliest work on KL-DRO dates back to 2013 by Hu & Hong (2012). However, it was only recently that these works have become widespread in deep learning. The RECOVER algorithm by Qi et al. (2021) was one of the early works to scale KL-DRO to deep neural networks. It attempted to solve the non-convex DRO problem with a duality-free stochastic method by formulating the min-max formulation into an equivalent stochastic compositional problem. ABSGD (Qi et al., 2023b) and SCDRO (Qi et al., 2023a) improved upon RECOVER by designing more efficient algorithms. However, the performance of these algorithms on large-scale models and datasets is not rigorously studied, as the Imagenet-LT, iNaturalist experiments conducted in these works started from a pre-trained network and only finetuned the last layer. In contrast, in this work, we learn the entire ViT-S model from scratch and show improved generalization.
@@ Appendix B Proofs of Section 3
$\min_{\theta\in\Theta}\frac{1}{\gamma}\log{\mathbb{E}_{\widehat{P}_{\text{data}}}[e^{\gamma\ell(z;\theta)}]},$
@@
For any fixed $\beta,\theta$, the inner supremum is attained at a $P^{\prime}$ that satisfies (see Theorem 1 of Hsieh et al. (2019))